A tree if less than three feet high and when transplanted has a 40% chance of survival. If six such trees are transplanted, what is the probability that at most four five will not survive?


1
Expert's answer
2022-05-26T11:20:04-0400

Let "X=" the number of trees which will not survive: "X\sim Bin (n, p)."

Given "n=6, p=1-0.4=0.6, q=0.4."


a.

"P(X\le4)=1-P(X=5)-P(X=6)"

"=1-\dbinom{6}{5}(0.6)^{5}(0.4)^{6-5}-\dbinom{6}{6}(0.6)^{6}(0.4)^{6-6}"

"=0.54432"

b.

"P(X\le5)=1-P(X=6)"

"=1-\dbinom{6}{6}(0.6)^{6}(0.4)^{6-6}=0.953344"